Consider the following reaction at 600K: A reacti
Answer #1

i) the equilibrium constant is given by

K_{p}=\frac{\left [P_{ HI} \right ]^{2}}{\left [P_{ H_{2} }\right ]\left [P_{ I_{2}} \right ]}

K_{p}=\frac{\left [1.26 \right ]^{2}}{\left [0.17\right ]\left [0.12\right ]}= 77.82

\therefore K_{p}= 77.82

ii) for the second reaction mixture

Reaction\: Quotient \: Q=\frac{\left [P_{ HI} \right ]^{2}}{\left [P_{ H_{2} }\right ]\left [P_{ I_{2}} \right ]}

Q=\frac{\left [1.1 \right ]^{2}}{\left [0.135\right ]\left [0.135 \right ]}=66.39

Since Q<K the reaction is not equilibrium.